The gantry crane of Prob 2/105 is repeated here, and the tension in cable AB is 14kN. Replace this force as it acts on point A by an equivalent force couple system at O. Point B is located at the center of the container top.


1
Expert's answer
2021-10-24T18:20:37-0400

"F_1=F_{AB}\cos \frac {20+3}{16+8+9+3}=11.2~N,"

"F_2=F_{AB}\sin\frac{5+9+3}{16+8+9+3}=6 .4~N."